OpenStudy (anonymous):

What is the slope and y-intercept of the equation 3(y − 2) + 6(x + 1) − 2 = 0?

OpenStudy (anonymous):

= 3y-6+6x+6-2=0 => 3y=2-6x so y=2/3 - 2x; the coefficient of x is the slope, which is -2. the y-intercept is when you plug 0 into x, then 3y=2-6*0 => y-intercept=2/3